OpenStudy (anonymous):

When solving a radical equation, Beth and Kelly came to two different conclusions. Beth found a solution, while Kelly's solution did not work in the equation. Create and justify two situations: one situation where Beth is correct and a separate situation where Kelly is correct.

OpenStudy (solomonzelman):

you need 2 radical equations. one with extraneous solution, and the other with a fully working solution (the one that works for checking).

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

I am not very good at coming up with these examples, but you can get ideas online....

OpenStudy (solomonzelman):

I found \(\Large\color{black}{ \sqrt{x-1}=x-7}\), but there are 2 solutions, and one works and one doesn't. You apparently, though, need a situation with all extraneous or with 1 extraneous solution, and a situation with all working or with 1 working solution.